OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 5 interactions. If there are 4 children, there can be 14 interactions. Which recursive equation represents the pattern? an = an – 1 + 2(n – 1) an = an – 1 + (n – 1) 2 an = an – 1 + 2(n – 1) an = an – 1 + (2n – 1)

OpenStudy (anonymous):

\[an = an-1 + (n-1)^{2}\] 14 = 5 + (4-1)^2 5 = 1 + (3-1)^2